OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax and x represents the original price of the item. : C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol. Numerical Answers Expected!

OpenStudy (anonymous):

store offers 15% off employees get an additional 10% whats chris's discount

OpenStudy (anonymous):

20 %

OpenStudy (anonymous):

now 15%+10%

OpenStudy (anonymous):

no*

OpenStudy (anonymous):

C = 0.765x + 0.06(0.765x) C=.765*100+.0435*100 C=76.50+4.35 C=$80.85 IS THE DISCOUNTED PRICE.

OpenStudy (anonymous):

so it is 80.85

OpenStudy (anonymous):

@A.K.C.15 let him do some of the work he needs to understand how to do it himself

OpenStudy (anonymous):

yes but look at the way to solve not just the answer

OpenStudy (anonymous):

thanks you helped alot

OpenStudy (anonymous):

do you understand how to solve for it ?@bubbasangster